# Optimal preconditioning techniques for finite volume approximation of three-dimensional conservative space-fractional diffusion equations

### v1

@@ 1 Introduction
Fractional diffusion equations (FDEs), as generalizations of integer-order diffusion equations, have been proven to be effective models for anomalous diffusion BWM1, BWM2, CLZ1, biology Magin1, and finance raberto2002waiting. Conventionally, spatial FDEs are formed by replacing second-order derivatives in classic diffusion equations with fractional Riemann-Liouville (R-L) or Riesz derivatives.
@@
Similar to their integer-order counterparts, FDEs rarely yield analytical solutions, making numerical approaches essential. In recent years, significant progresses have been made in this area, including finite difference (FD) methods CD2, LL1, lin2021stability, she2022class, SL1, TZD1, ZD1, finite element methods bu2014galerkin, JLPZ1, zhao2015finite, and finite volume (FV) methods donatelli2018spectral, fang2022fast, feng2015stability, fu2019finite, fu2019stability, hejazi2014stability, kong2024fast, LZTBA1, liu2021analysis, pan2017fast, pan2016fast, wang2013superfast, simmons2017finite, zheng2019efficient.
However, unlike classical diffusion equations, the nonlocal nature of fractional derivatives typically yields full and dense discrete linear systems that require $\mathcal{O}(N^{2})$ memory storage and $\mathcal{O}(N^{3})$ computational cost for the matrix size $N$. Consequently, the computational cost for large-scale simulations of FDEs become prohibitively expensive. Notably, the discrete linear systems derived from the FD and FV methods exhibit Toeplitz-like structures. By leveraging these structures, the matrix-vector multiplications in the Krylov subspace method can be efficiently computed using fast Fourier transforms (FFTs). However, numerical experiments demonstrate that the ill-conditioning of the coefficient matrix increasingly degrades the performance of the iterative solvers as the matrix size $N$ grows. In such cases, preconditioning techniques are employed to accelerate the convergence of Krylov subspace methods. Classical preconditioners, including circulant preconditioners lei2016multilevel, lei2013circulant, banded preconditioners JLZ1, LYJ1, SLYL1, approximate inverse preconditioners PKNS1, zeng2022tau, structure preserving preconditioners DMS1, and $\tau$-preconditioners huang2021spectral, huang2023tau, lin2023tau, qu2025novel, have been proposed to accelerate the solution of the discrete linear systems arising from FD discretization.
@@
In early work, Wang and Du wang2013superfast have studied a superfast-preconditioned conjugate gradient squared method for one-dimensional (1D) steady-state Riesz space FDEs. Their method reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$ per iteration and reduces the memory requirement from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. Subsequently, Pan et al. pan2017fast, pan2016fast extended this problem to R-L space FDEs in two dimensions, including both time-independent and time-dependent variable-coefficient cases. They respectively proposed a scaled-circulant preconditioner and an approximate inverse-circulant preconditioner for solving the linear systems arising from the FV discretization. Theoretical analyses demonstrate that for the time-independent case, the difference between the scaled-circulant preconditioner and the coefficient matrix equals the sum of a small-norm matrix and a low-rank matrix, while for the time-dependent problem, the spectra of the corresponding preconditioned matrices are clustered around $1$. Then rapid convergence of the preconditioned Krylov subspace methods can be expected with these properties. In donatelli2018spectral, Donatelli et al. developed a finite volume discretization for 1D and two-dimensional (2D) conservative steady-state FDEs with variable coefficients. They revealed the intrinsic nature of the resulting sequence of coefficient matrices through the framework of generalized locally Toeplitz sequences. This analysis provides essential spectral information for designing efficient preconditioners and multigrid methods to solve the resulting linear systems. Subsequently, Fu’s group fu2019finite, fu2019stability combined Crank-Nicolson (CN) and FV methods for solving 1D and 2D conservative SFDEs. They established the stability and convergence of the CN-FV scheme through rigorous norm analysis. To improve computational efficiency, circulant preconditioners were designed to solve the resulting linear systems, although the spectral analysis of these preconditioned matrices was not addressed. Additionally, the CN-FV scheme has been effectively extended to 3D nonlinear distributed-order conservative SFDEs, see zheng2019efficient. Furthermore, to handle high-dimensional conservative FDEs, the alternating direction implicit method can be employed. For more details, readers are referred to chou2021finite, liu2021analysis and the references therein.
@@
The aforementioned works demonstrate the strong potential of the fast algorithms for conservative SFDEs (1.1). To the best of our knowledge, there is still a lack of sine transform-based preconditioning techniques for conservative SFDEs. In this paper, based on the CN-FV scheme developed by Fu’s group fu2019finite, zheng2019efficient, we propose a novel sine transform-based preconditioning strategy combined with the CG and GMRES methods to solve the symmetric and non-symmetric discrete linear systems arising from conservative SFDEs. The aim of this paper is twofold:
@@ 2 Crank-Nicolson and finite volume approximations
For simplicity, we denote $\delta_{i}=\begin{cases}\alpha,&i=1\\ \beta,&i=2\\ \gamma,&i=3\end{cases},$ then the CN-FV scheme (2) can be reformulated in the matrix form given in fu2019finite, zheng2019efficient
@@
As discussed in fu2019finite, zheng2019efficient, the scheme in (2.10) can be shown to be unconditionally stable and convergent, achieving second-order accuracy in time and $\min\{1+\alpha,1+\beta,1+\gamma\}$ accuracy in space with respect to a weighted discrete norm. In particular, when $k_{i,+}=k_{i,-}\,(i=1,2,3)$, (1.1) reduces to a symmetry problem, then the spatial accuracy of the scheme (2.10) can be improved to second-order, as detailed in fu2019stability.
@@
fu2019stability, qu2021fast For all $0<\delta_{i}<1$, and let $\{q_{k}^{(\delta_{i})}\}_{k=0}^{n_{i}}$ be defined as (2.14). Then we have
@@ 3 Preconditioners and convergence behaviors
In general, the iterative method converges slowly since the coefficient matrix will be ill-conditioned as the spatial grid mesh increases. Hence, some specific techniques, such as preconditioning, are necessary to improve the proposed fast iterative method by reducing the iteration number. Recently, sine transform-based preconditioners have been widely used to precondition both symmetric and non-symmetric Toeplitz systems derived from FD methods, as illustrated in hon2024symbol, huang2021spectral, huang2024optimal, huang2023tau, li2025multilevel, lin2023tau, qin2023sine, qu2025novel, zhang2022fast and the references therein. Inspired by this novel preconditioning idea, we aim at constructing the following preconditioner for solving the CN-FV scheme (2.16):
@@
where $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ is the $\tau$ preconditioner bini1990new for $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$. Note that $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ can be fast diagonalized by the sine transform matrix $\mathbf{S}_{n_{i}}$ with $\mathbf{S}_{n_{i}}]_{j,k}=\sqrt{\frac{2}{n_{i}+1}}\sin\left(\frac{jk\pi}{n_{i}+1}\right)$ for $1\leq j,k\leq n_{i}$; and the eigenvalues of $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ can be determined by the first column elements of $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ bini1990new, huang2023tau. Since the symmetric tridiagonal matrix $\mathbf{A}_{n_{i}}$ is also a $\tau$ matrix, the preconditioner $\mathbf{P}$ can be diagonalized by the 3-level sine transform matrix $\mathbf{S}_{N}$, i.e.,
@@
The results can be obtained directly by using the Geršgorin disk theorem bai2021matrix, axelsson1996iterative, varga2011gervsgorin. ∎
@@
In fact, it has been established in fu2019stability that $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ is SPD. In the following, we show that matrix $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ is SPD as well. Let $\lambda_{k}$ be the $k$-th eigenvalue of $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ for $k=1,2,\ldots,n_{i}$, then $\lambda_{k}$ can be determined by using the first column elements of $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$, i.e.,
@@
From (2.12), we know that when $k_{i,+}=k_{i,-}\,(i=1,2,3)$, $\mathbf{B}_{n_{i}}^{\delta_{i}}$ defined in (2.12) is SPD fu2019stability, which combined with Lemma 3.1 indicates that the linear systems (2.10) are SPD. In this case, the preconditioned CG (PCG) method with the preconditioner $\mathbf{P}$ is used to solve the systems (2.10). Otherwise, when $k_{1,+}\not=k_{1,-}$ or $k_{2,+}\not=k_{2,-}$ or $k_{3,+}\not=k_{3,-}$, the matrix $\mathbf{B}_{n_{i}}^{\delta_{i}}$ is non-symmetric, then the PGMRES method with the preconditioner $\mathbf{P}$ will be adopted to solve (2.10). In what follows, the convergence behaviors of the two methods will be discussed in detail based on the following important inequality.
@@
lin2018efficient For nonnegative numbers $\xi_{i}$ and positive numbers $\zeta_{i}$ $(1\leq i\leq m)$, it holds that
@@ 3.1 Convergence of the CG method for symmetric systems
It is well known that the convergence rate of the CG method can be described by the eigenvalue distribution of the coefficient matrix. The more clustered the eigenvalues are, the faster the convergence rate will be. In particular, if the eigenvalues of the preconditioned matrix are bounded with mesh-size independent bounds, then the CG method converges linearly chan2007introduction, huang2021spectral. In this subsection, we will show the effectiveness of the PCG method for (2.10) in the symmetric case by showing the bounded eigenvalue distribution of the preconditioned matrix $\mathbf{P}^{-1/2}\mathbf{A}\mathbf{P}^{-1/2}$.
@@
huang2021spectral Let $T_{n}=\left[t_{|i-j|}\right]$ be a symmetric Toeplitz matrix and $\tau(T_{n})$ be the corresponding $\tau$ matrix. If the entries of $T_{n}$ are equipped with the following properties,
@@
where $\mathbf{A}_{n_{i}}=\mathbf{L}_{n_{i}}\mathbf{L}_{n_{i}}^{T}$ for $i=2,3$ is the Cholesky factorization of the SPD matrix $\mathbf{A}_{n_{i}}$. By (3.7), the Rayleigh quotient theorem chen2005matrix and properties of the Kronecker product, we have
@@ 3.2 Convergence of the GMRES method for non-symmetric systems
Unlike the CG method, the behavior of the GMRES cannot be determined from eigenvalues alone greenbaum1997iterative. However, residual-norm estimates provide practical ways to characterize its convergence; see, e.g., lin2023tau, lin2021all, liu2020fast, qin2023sine.
@@
[elman2014finite, Proposition 7.3] Let $\mathbf{Z}\mathbf{v}=\mathbf{w}$ be a real square linear system with $\mathcal{H}(\mathbf{Z})\succ\mathcal{O}$. Then, the residuals of the iterates generated by applying (restarted or non-restarted) GMRES method with an arbitrary initial guess to solve $\mathbf{Z}\mathbf{v}=\mathbf{w}$ satisfy
@@
which, combined with the Rayleigh quotient theorem chen2005matrix for real symmetric matrix implies
@@
For $\delta_{i}\in\mathbbm{C}$ with $\textrm{Re}(\delta_{i})>0$, it has been shown that $|q_{i}^{(\delta_{i})}|=\mathcal{O}(i^{-1-\textrm{Re}(\delta_{i})})$, see Lemma 3.4 in donatelli2018spectral, then $g_{\delta_{i}}(\theta)$ is analytic on $\{s\in\mathbbm{C}\big|\textrm{Re}(s)>0\}$ with respect to $\delta_{i}$. Hence with the analytic continuation of the Lerch transcendent bateman1953higher $\Phi(z,s,\nu)=\sum_{k=0}^{\infty}(k+\nu)^{-s}z^{k}$ that is
@@
From the proof of Lemma 3.8, we note that the explicit expression of generating function $g_{\delta_{i}}(\theta)$ for the Toeplitz matrix $\mathbf{T}_{n_{i}}^{\delta_{i}}$ agrees with the result in chou2021finite. However, the proofs of two results are quite different, and our method appears to be simpler. Moreover, Lemma 3.8 also provides the results for $\operatorname{Re}(g_{\delta_{i}}(\theta))$, $\operatorname{Im}(g_{\delta_{i}}(\theta))$ and $\frac{\left|\operatorname{Im}(g_{\delta_{i}}(\theta))\right|}{\operatorname{Re}(g_{\delta_{i}}(\theta))}$, which are not included in chou2021finite.
Based on Lemma 3.8, and by applying the same argument as in Corollary 3.1 of lin2023tau, we can readily derive the following lemma.
@@
As has been pointed out in fu2019stability that $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ is SPD. By Lemmas 3.1 and 3.3, it follows that $\mathcal{H}\left(\mathbf{P}^{-\frac{1}{2}}\mathbf{A}\mathbf{P}^{-\frac{1}{2}}\right)$ is positive definite. Therefore, Lemma 3.6 can be used.
@@
Noting that $\lambda_{\min}(\mathbf{P})\geq\lambda_{\min}(\mathbf{A}_{N})\geq(\frac{1}{2})^{3}=\frac{1}{8}$ from Lemma 3.1, the following theorem follows directly from the proofs in huang2024optimal, huang2025efficient, lin2024single.

### v2

@@ 1 Introduction
Fractional diffusion equations (FDEs), as generalizations of integer-order diffusion equations, have been proven to be effective models for anomalous diffusion Benson et al. [2000a, b], Carreras et al. [2001], biology Magin [2006], and finance Raberto et al. [2002]. Conventionally, spatial FDEs are formed by replacing second-order derivatives in classic diffusion equations with fractional Riemann-Liouville (R-L) or Riesz derivatives.
@@
Similar to their integer-order counterparts, FDEs rarely yield analytical solutions, making numerical approaches essential. In recent years, significant progresses have been made in this area, including finite difference (FD) methods Çelik and Duman [2012], Lin and Liu [2020], Lin and She [2021], She [2022], Sousa and Li [2015], Tian et al. [2015], Zhao and Deng [2016], finite element methods Bu et al. [2014], Jin et al. [2014], Zhao et al. [2015], and finite volume (FV) methods Donatelli et al. [2018], Fang et al. [2022], Feng et al. [2015], Fu et al. [2019a, b], Hejazi et al. [2014], Kong and Fu [2024], Liu et al. [2014, 2021], Pan et al. [2017, 2016], Wang and Du [2013], Simmons et al. [2017], Zheng et al. [2019].
However, unlike classical diffusion equations, the nonlocal nature of fractional derivatives typically yields full and dense discrete linear systems that require $\mathcal{O}(N^{2})$ memory storage and $\mathcal{O}(N^{3})$ computational cost for the matrix size $N$. Consequently, the computational cost for large-scale simulations of FDEs become prohibitively expensive. Notably, the discrete linear systems derived from the FD and FV methods exhibit Toeplitz-like structures. By leveraging these structures, the matrix-vector multiplications in the Krylov subspace method can be efficiently computed using fast Fourier transforms (FFTs). However, numerical experiments demonstrate that the ill-conditioning of the coefficient matrix increasingly degrades the performance of the iterative solvers as the matrix size $N$ grows. In such cases, preconditioning techniques are employed to accelerate the convergence of Krylov subspace methods. Classical preconditioners, including circulant preconditioners Lei et al. [2016], Lei and Sun [2013], banded preconditioners Jin et al. [2015], Lin et al. [2014], She et al. [2021], approximate inverse preconditioners Pan et al. [2014], Zeng et al. [2022], structure preserving preconditioners Donatelli et al. [2016], and $\tau$-preconditioners Huang et al. [2022, 2023], Lin et al. [2023], Qu et al. [2025], have been proposed to accelerate the solution of the discrete linear systems arising from FD discretization.
@@
In early work, Wang and Du Wang and Du [2013] have studied a superfast-preconditioned conjugate gradient squared method for one-dimensional (1D) steady-state Riesz space FDEs. Their method reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$ per iteration and reduces the memory requirement from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. Subsequently, Pan et al. Pan et al. [2017, 2016] extended this problem to R-L space FDEs in two dimensions, including both time-independent and time-dependent variable-coefficient cases. They respectively proposed a scaled-circulant preconditioner and an approximate inverse-circulant preconditioner for solving the linear systems arising from the FV discretization. Theoretical analyses demonstrate that for the time-independent case, the difference between the scaled-circulant preconditioner and the coefficient matrix equals the sum of a small-norm matrix and a low-rank matrix, while for the time-dependent problem, the spectra of the corresponding preconditioned matrices are clustered around $1$. Then rapid convergence of the preconditioned Krylov subspace methods can be expected with these properties. In Donatelli et al. [2018], Donatelli et al. developed a finite volume discretization for 1D and two-dimensional (2D) conservative steady-state FDEs with variable coefficients. They revealed the intrinsic nature of the resulting sequence of coefficient matrices through the framework of generalized locally Toeplitz sequences. This analysis provides essential spectral information for designing efficient preconditioners and multigrid methods to solve the resulting linear systems. Subsequently, Fu’s group Fu et al. [2019a, b] combined Crank-Nicolson (CN) and FV methods for solving 1D and 2D conservative SFDEs. They established the stability and convergence of the CN-FV scheme through rigorous norm analysis. To improve computational efficiency, circulant preconditioners were designed to solve the resulting linear systems, although the spectral analysis of these preconditioned matrices was not addressed. Additionally, the CN-FV scheme has been effectively extended to 3D nonlinear distributed-order conservative SFDEs, see Zheng et al. [2019]. Furthermore, to handle high-dimensional conservative FDEs, the alternating direction implicit method can be employed. For more details, readers are referred to Chou and Lei [2021], Liu et al. [2021] and the references therein.
@@
The aforementioned works demonstrate the strong potential of the fast algorithms for conservative SFDEs (1.1). To the best of our knowledge, there is still a lack of sine transform-based preconditioning techniques for conservative SFDEs. In this paper, based on the CN-FV scheme developed by Fu’s group Fu et al. [2019a], Zheng et al. [2019], we propose a novel sine transform-based preconditioning strategy combined with the CG and GMRES methods to solve the symmetric and non-symmetric discrete linear systems arising from conservative SFDEs. The aim of this paper is twofold:
@@ 2 Crank-Nicolson and finite volume approximations
For simplicity, we denote $\delta_{i}=\begin{cases}\alpha,&i=1\\ \beta,&i=2\\ \gamma,&i=3\end{cases},$ then the CN-FV scheme (2) can be reformulated in the matrix form given in Fu et al. [2019a], Zheng et al. [2019]
@@
As discussed in Fu et al. [2019a], Zheng et al. [2019], the scheme in (2.10) can be shown to be unconditionally stable and convergent, achieving second-order accuracy in time and $\min\{1+\alpha,1+\beta,1+\gamma\}$ accuracy in space with respect to a weighted discrete norm. In particular, when $k_{i,+}=k_{i,-}\,(i=1,2,3)$, (1.1) reduces to a symmetry problem, then the spatial accuracy of the scheme (2.10) can be improved to second-order, as detailed in Fu et al. [2019b].
@@
Fu et al. [2019b], Qu and Li [2021] For all $0<\delta_{i}<1$, and let $\{q_{k}^{(\delta_{i})}\}_{k=0}^{n_{i}}$ be defined as (2.14). Then we have
@@ 3 Preconditioners and convergence behaviors
In general, the iterative method converges slowly since the coefficient matrix will be ill-conditioned as the spatial grid mesh increases. Hence, some specific techniques, such as preconditioning, are necessary to improve the proposed fast iterative method by reducing the iteration number. Recently, sine transform-based preconditioners have been widely used to precondition both symmetric and non-symmetric Toeplitz systems derived from FD methods, as illustrated in Hon et al. [2025], Huang et al. [2022, 2025b, 2023], Li and Hon [2025], Lin et al. [2023], Qin et al. [2023], Qu et al. [2025], Zhang et al. [2022] and the references therein. Inspired by this novel preconditioning idea, we aim at constructing the following preconditioner for solving the CN-FV scheme (2.16):
@@
where $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ is the $\tau$ preconditioner Bini and Benedetto [1990] for $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$. Note that $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ can be fast diagonalized by the sine transform matrix $\mathbf{S}_{n_{i}}$ with $\mathbf{S}_{n_{i}}]_{j,k}=\sqrt{\frac{2}{n_{i}+1}}\sin\left(\frac{jk\pi}{n_{i}+1}\right)$ for $1\leq j,k\leq n_{i}$; and the eigenvalues of $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ can be determined by the first column elements of $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ Bini and Benedetto [1990], Huang et al. [2023]. Since the symmetric tridiagonal matrix $\mathbf{A}_{n_{i}}$ is also a $\tau$ matrix, the preconditioner $\mathbf{P}$ can be diagonalized by the 3-level sine transform matrix $\mathbf{S}_{N}$, i.e.,
@@
The results can be obtained directly by using the Geršgorin disk theorem Bai and Pan [2021], Axelsson [1996], Varga [2011]. ∎
@@
In fact, it has been established in Fu et al. [2019b] that $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ is SPD. In the following, we show that matrix $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ is SPD as well. Let $\lambda_{k}$ be the $k$-th eigenvalue of $\tau(\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}}))$ for $k=1,2,\ldots,n_{i}$, then $\lambda_{k}$ can be determined by using the first column elements of $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$, i.e.,
@@
From (2.12), we know that when $k_{i,+}=k_{i,-}\,(i=1,2,3)$, $\mathbf{B}_{n_{i}}^{\delta_{i}}$ defined in (2.12) is SPD Fu et al. [2019b], which combined with Lemma 3.1 indicates that the linear systems (2.10) are SPD. In this case, the preconditioned CG (PCG) method with the preconditioner $\mathbf{P}$ is used to solve the systems (2.10). Otherwise, when $k_{1,+}\not=k_{1,-}$ or $k_{2,+}\not=k_{2,-}$ or $k_{3,+}\not=k_{3,-}$, the matrix $\mathbf{B}_{n_{i}}^{\delta_{i}}$ is non-symmetric, then the PGMRES method with the preconditioner $\mathbf{P}$ will be adopted to solve (2.10). In what follows, the convergence behaviors of the two methods will be discussed in detail based on the following important inequality.
@@
Lin et al. [2018] For nonnegative numbers $\xi_{i}$ and positive numbers $\zeta_{i}$ $(1\leq i\leq m)$, it holds that
@@ 3.1 Convergence of the CG method for symmetric systems
It is well known that the convergence rate of the CG method can be described by the eigenvalue distribution of the coefficient matrix. The more clustered the eigenvalues are, the faster the convergence rate will be. In particular, if the eigenvalues of the preconditioned matrix are bounded with mesh-size independent bounds, then the CG method converges linearly Chan and Jin [2007], Huang et al. [2022]. In this subsection, we will show the effectiveness of the PCG method for (2.10) in the symmetric case by showing the bounded eigenvalue distribution of the preconditioned matrix $\mathbf{P}^{-1/2}\mathbf{A}\mathbf{P}^{-1/2}$.
@@
Huang et al. [2022] Let $T_{n}=\left[t_{|i-j|}\right]$ be a symmetric Toeplitz matrix and $\tau(T_{n})$ be the corresponding $\tau$ matrix. If the entries of $T_{n}$ are equipped with the following properties,
@@
where $\mathbf{A}_{n_{i}}=\mathbf{L}_{n_{i}}\mathbf{L}_{n_{i}}^{T}$ for $i=2,3$ is the Cholesky factorization of the SPD matrix $\mathbf{A}_{n_{i}}$. By (3.7), the Rayleigh quotient theorem Chen [2005] and properties of the Kronecker product, we have
@@ 3.2 Convergence of the GMRES method for non-symmetric systems
Unlike the CG method, the behavior of the GMRES cannot be determined from eigenvalues alone Greenbaum [1997]. However, residual-norm estimates provide practical ways to characterize its convergence; see, e.g., Lin et al. [2023], Lin and Ng [2021], Liu and Wu [2020], Qin et al. [2023].
@@
[Elman et al., 2014, Proposition 7.3] Let $\mathbf{Z}\mathbf{v}=\mathbf{w}$ be a real square linear system with $\mathcal{H}(\mathbf{Z})\succ\mathcal{O}$. Then, the residuals of the iterates generated by applying (restarted or non-restarted) GMRES method with an arbitrary initial guess to solve $\mathbf{Z}\mathbf{v}=\mathbf{w}$ satisfy
@@
which, combined with the Rayleigh quotient theorem Chen [2005] for real symmetric matrix implies
@@
For $\delta_{i}\in\mathbbm{C}$ with $\textrm{Re}(\delta_{i})>0$, it has been shown that $|q_{i}^{(\delta_{i})}|=\mathcal{O}(i^{-2-\textrm{Re}(\delta_{i})})$, see Lemma 3.4 in Donatelli et al. [2018], then $g_{\delta_{i}}(\theta)$ is analytic on $\{s\in\mathbbm{C}\big|\textrm{Re}(s)>0\}$ with respect to $\delta_{i}$. Hence with the analytic continuation of the Lerch transcendent Bateman [1953] $\Phi(z,s,\nu)=\sum_{k=0}^{\infty}(k+\nu)^{-s}z^{k}$ that is
@@
From the proof of Lemma 3.8, we note that the explicit expression of generating function $g_{\delta_{i}}(\theta)$ for the Toeplitz matrix $\mathbf{T}_{n_{i}}^{\delta_{i}}$ agrees with the result in Chou and Lei [2021]. However, the proofs of two results are quite different, and our method appears to be simpler. Moreover, Lemma 3.8 also provides the results for $\operatorname{Re}(g_{\delta_{i}}(\theta))$, $\operatorname{Im}(g_{\delta_{i}}(\theta))$ and $\frac{\left|\operatorname{Im}(g_{\delta_{i}}(\theta))\right|}{\operatorname{Re}(g_{\delta_{i}}(\theta))}$, which are not included in Chou and Lei [2021].
Based on Lemma 3.8, and by applying the same argument as in Corollary 3.1 of Lin et al. [2023], we can readily derive the following lemma.
@@
As has been pointed out in Fu et al. [2019b] that $\mathcal{H}(\mathbf{T}_{n_{i}}^{\delta_{i}})$ is SPD. By Lemmas 3.1 and 3.3, it follows that $\mathcal{H}\left(\mathbf{P}^{-\frac{1}{2}}\mathbf{A}\mathbf{P}^{-\frac{1}{2}}\right)$ is positive definite. Therefore, Lemma 3.6 can be used.
@@
Noting that $\lambda_{\min}(\mathbf{P})\geq\lambda_{\min}(\mathbf{A}_{N})\geq(\frac{1}{2})^{3}=\frac{1}{8}$ from Lemma 3.1, the following theorem follows directly from the proofs in Huang et al. [2025b, a], Lin et al. [2024].